# Abelian congruences and similarity in varieties with a weak difference term

### v1

@@ 1. Introduction
Arguably the most important result in universal algebra in the last ten years is the positive resolution to the Constraint Satisfaction Problem (CSP) Dichotomy Conjecture, announced independently in 2017 by Andrei Bulatov [2] and Dmitriy Zhuk [36, 37]. One particular feature of Zhuk’s proof is his analysis of “rectangular critical” subdirect products $R\leq_{sd}{\mathbf{A}}_{1}\times\cdots\times{\mathbf{A}}_{n}$ of finite algebras ${\mathbf{A}}_{1},\ldots,{\mathbf{A}}_{n}$ in certain locally finite idempotent Taylor varieties. Zhuk showed that, in the varieties he considers, such subdirect products generate derived relations, which he called “bridges,” between certain meet-irreducible congruences $\rho_{i}\in\operatorname{Con}({\mathbf{A}}_{i})$ which $R$ determines. Zhuk established a number of useful properties of his bridges and used them to tease out implicit linear equations in the CSP instances he considered.
@@
Strikingly, Zhuk’s definition of “bridge” formally resembles some aspects of Hagemann and Herrmann’s formulation [11] from 1979 of the centrality relation on congruences of a single algebra. Moreover, the “there-exists-a-bridge” relation, under suitable restrictions, has an uncanny resemblance to the “similarity” relation of Freese [5] from 1983; this latter relation is between subdirectly irreducible algebras in a congruence modular variety, and can be viewed as an amplification of a certain correspondence involving the centrality relation. As well, Zhuk’s bridges arising in rectangular critical subdirect products in locally finite idempotent Taylor varieties resemble the “graphs of similarity” which Kearnes and Szendrei proved arise in rectangular critical relations of the form $R\leq{\mathbf{A}}^{n}$ in congruence modular varieties [20, Theorem 2.5(6)].
@@
Our main goal in this paper and two companion papers [33, 32] is to establish precise connections between Zhuk’s bridges, centrality, and similarity, and then to apply these connections to the study of rectangular critical relations. In order to connect Zhuk’s bridges to similarity, we need to first generalize the similarity relation from its original setting (congruence modular varieties) to a broader class also containing all locally finite idempotent Taylor varieties. For this purpose, we focus on the class of varieties “having a weak difference term.” Here a “weak difference term” is an idempotent ternary term that specializes to a Maltsev operation whenever it is restricted to a block of an abelian congruence. Such terms are weak versions of terms, known as “difference terms,” that exist and play an important role in congruence modular varieties [9]. Weak difference terms arose organically in tame congruence theory in the 1980s, when Hobby and McKenzie [13] showed that every locally finite Taylor has such a term. The property of a (general) variety having a weak difference term was first explicitly identified in the 1990s in work of Lipparini [23], Kearnes [15], and Kearnes and Szendrei [19], who explored the extent to which centralizer properties in congruence modular varieties extend to varieties with a weak difference term. Recent work concerning varieties with a weak difference term is represented by [17, 35, 21, 16].
@@
In this paper we show that Freese’s similarity relation generalizes nicely to varieties with a weak difference term. It turns out that in the abelian monolith case, Freese’s original definition of similarity, as well as the reformulation in [6], does not generalize in a straightforward way. Instead, we take as our starting point two equivalent (in the congruence modular setting) characterizations of similarity due to Freese – one based on the construction from Hagemann and Herrmann’s seminal work, the other a congruence lattice-theoretic formulation – and show that both characterizations generalize to, and are provably equivalent in, varieties with a weak difference term (Theorems 4.6 and 7.8). We also give a third equivalent characterization using a variant of Zhuk’s bridges, which we call “proper bridges” (Definition 7.3).
Along the way, we explore the abelian groups arising from the congruence classes of an abelian congruence $\theta$ in an algebra in a variety having a weak difference term. While the situation is not as well-behaved as in the congruence modular case (see Example 8.4), we discern some structure in Section 5. In particular, we show that the construction of Hagemann and Herrmann mentioned above provides a sort of “universal domain” for the abelian groups arising from the $\theta$-classes of ${\mathbf{A}}$ within each $(0:\theta)$-class, where $(0:\theta)$ is the largest congruence centralizing $\theta$ (Lemmas 5.2 and 5.4(3) and Corollary 5.8).
@@
We also expend some energy explaining a folklore result: that each class of an abelian congruence $\theta$ (in an algebra ${\mathbf{A}}$ with a weak difference term) naturally supports the structure of a left vector space over a division ring $\mathbb{F}$, provided $\theta$ is minimal. Here $\mathbb{F}$ depends on ${\mathbf{A}}$ and $\theta$ but not on the particular $\theta$-class. In particular, we show that if two subdirectly irreducible algebras in a variety with a weak difference term have abelian monoliths and are similar, then the division rings determined by their monoliths are isomorphic.
@@
In the companion paper [33], we will develop Zhuk’s theory of bridges (in a slightly more general setting than Zhuk did), and describe the precise relationships between the objects in Zhuk’s theory, the centrality relation, similarity, and proper bridges as defined here. In a third paper [32], we will study rectangular critical relations in locally finite Taylor varieties, extending the work of Zhuk [37] and connecting these relations to similarity (as was done by Kearnes and Szendrei [20] in the congruence modular setting). In particular, we will show explicitly how each such relation can be understood as encoding a family of “parallel” linear equations over a finite field.
Acknowledgment. We thank Keith Kearnes for several helpful conversations, including one which led us to a proof of an early version of Theorem 4.6.
@@ 2. Definitions
If ${\mathbf{A}}$ is an algebra and $\alpha,\beta\in\operatorname{Con}({\mathbf{A}})$ with $\alpha\leq\beta$, then $\beta/\alpha$ denotes the congruence of ${\mathbf{A}}/\alpha$ corresponding to $\beta$ via the Correspondence Theorem ([26, Theorem 4.12] or [1, Theorem 3.6]). If in addition $\gamma,\delta\in\operatorname{Con}({\mathbf{A}})$ with $\gamma\leq\delta$, then we write $(\alpha,\beta)\nearrow(\gamma,\delta)$ and say that $(\alpha,\beta)$ is perspective up to $(\gamma,\delta)$, or transposes up to $(\gamma,\delta)$, if $\beta\wedge\gamma=\alpha$ and $\beta\vee\gamma=\delta$. The notation $(\gamma,\delta)\searrow(\alpha,\beta)$ means the same thing. We write $\alpha\prec\beta$ and say that $\beta$ covers $\alpha$, or is an upper cover of $\alpha$, if $\alpha<\beta$ and there does not exist a congruence $\gamma$ satisfying $\alpha<\gamma<\beta$. A congruence $\alpha$ is minimal if it covers 00, and is completely meet-irreducible if $\alpha\neq 1$ and there exists $\alpha^{+}$ with $\alpha\prec\alpha^{+}$ and such that $\alpha<\beta\implies\alpha^{+}\leq\beta$ for all $\beta\in\operatorname{Con}({\mathbf{A}})$. A subset $T\subseteq A$ is a transversal for a congruence $\alpha$ if it contains exactly one element from each $\alpha$-class.
@@
More generally, if $S,T$ are tolerances of ${\mathbf{A}}$ (i.e., reflexive symmetric subuniverses of ${\mathbf{A}}^{2}$), then we can define ${\mathbf{A}}(S,T)$, $X(S,T)$, $M(S,T)$, and the relation $C(S,T;\delta)$ in the same way; see [17, §2.5].
Let $\theta,\delta$ be congruences of an algebra ${\mathbf{A}}$. The centralizer (or annihilator) of $\theta$ modulo $\delta$, denoted $(\delta:\theta)$, is the unique largest congruence $\varphi$ for which $C(\varphi,\theta;\delta)$ holds. We say that $\theta$ is abelian if $C(\theta,\theta;0)$ holds; equivalently, if $\theta\leq(0:\theta)$. More generally, if $\theta,\delta\in\operatorname{Con}({\mathbf{A}})$ with $\delta\leq\theta$, then we say that $\theta$ is abelian modulo $\delta$ if any of the following equivalent conditions hold: (i) $\theta/\delta$ is an abelian congruence of ${\mathbf{A}}/\delta$; (ii) $C(\theta,\theta;\delta)$ holds; (iii) $\theta\leq(\delta:\theta)$. The following lemma expresses that centralizers behave nicely under certain quotients and surjective homomorphisms.
@@
At the level of varieties, the property of having a weak difference term can be characterized by the existence of idempotent terms satisfying a certain pattern of identities, i.e., by an idempotent Maltsev condition, as shown by Kearnes and Szendrei [19, Theorem 4.8]. (See Theorem 8.1 for a slight improvement.) In the context of locally finite varieties, the situation is simpler.
@@
This follows by combining [13, Theorem 9.6], [30, Corollary 5.3], [24, Theorem 2.2], and [19, Theorem 4.8]. ∎
Lemma 5.2 and Corollary 5.3 mention difference terms; these are weak difference terms which moreover satisfy the identity $d(x,x,y)\approx y$. By an old theorem of Gumm [9], every congruence modular variety has a difference term.
@@
if $a_{i}=b_{i}$ for $i=1,2,3$, then $a_{4}=b_{4}$. It is easy an easy exercise to show that if $\theta$ satisfies the two-term condition, then $\theta$ is abelian. The converse implication is also true if ${\mathbf{A}}$ belongs to a variety having a weak difference term, by [23, Theorem 3.1] or [19, Corollary 4.5]. (See Lemma 3.3 for a slight improvement.)
@@ 3. Tools
The first result is an easily proved generalization of the well-known fact (see e.g. [1, Theorem 4.65(1)]) that if an algebra ${\mathbf{A}}$ has a Maltsev term, then every reflexive subuniverse of ${\mathbf{A}}^{2}$ is a congruence of ${\mathbf{A}}$. (A proof can be easily extracted from the proof of [17, Corollary 6.22].)
@@
See Lipparini’s proof of [23, Theorem 3.1(a)]. His Theorem 3.1(a) assumes that ${\mathbf{A}}$ belongs to a variety with a difference term, but in fact all the argument requires is that ${\mathbf{A}}$ itself have a weak difference term. ∎
@@
Suppose ${\mathbf{A}}$ is an algebra having a weak difference term $d(x,y,z)$, and $\theta$ is an abelian congruence of ${\mathbf{A}}$. Then for all $e\in A$, the structure $(e/\theta,+,e)$ is an abelian group where $x+y:=d(x,e,y)$ for $x,y\in e/\theta$. Moreover, in this group we have $-x=d(e,x,e)$ and $d(x,y,z)=x-y+z$ for all $x,y,z\in e/\theta$.
@@
Let $B=e/\theta$. Then $(B,d{\upharpoonright}_{B^{3}})$ is an abelian algebra with Maltsev operation $d{\upharpoonright}_{B^{3}}$. Now apply Gumm [8, Theorem 4.7] or Herrmann’s main theorem from [12] (or see [31, Corollary 2] for a short proof). ∎
@@
Suppose ${\mathbf{A}}$ is an algebra and $\theta$ is an abelian congruence. If $d(x,y,z)$ and $t(x,y,z)$ are two weak difference terms for ${\mathbf{A}}$, then for every $\theta$-class $B$ we have that $d(a,b,c)=t(a,b,c)$ whenever $a\stackrel{{\scriptstyle\theta}}{{\equiv}}b\stackrel{{\scriptstyle\theta}}{{\equiv}}c$.
@@
Again this is a well-known consequence of $(B,d{\upharpoonright}_{B^{3}})$ being an abelian Maltsev operation. For $a,b,c\in B$, Proposition 3.4 yields
$d(t(a,b,b),t(b,b,b),t(b,b,c))=t(d(a,b,b),d(b,b,b),d(b,b,c))$
@@
The operation $+$ in the previous definition does not depend on the choice of the weak difference term, by Lemma 3.6. The notation $M(\theta,e)$ has been used in [6] and elsewhere for $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$. In case ${\mathbf{A}}$ is finite and $\theta$ is a minimal congruence, more can be said about the groups $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$. The following result is possibly folklore; we include it here for completeness.
@@
This was proved by Freese [5, §4] in the congruence modular case, and essentially the same proof works here. We recast his argument using tools from a recent paper of Mayr and Szendrei [25]. (See also the arguments in Section 6, where a generalization is proved.)
@@
First note that we can assume that ${\mathbf{A}}$ has no nullary operations (by replacing each nullary operation with a constant unary operation). Let ${\mathbf{I}}={\mathbf{A}}/\theta$, and let $\nu_{\theta}:{\mathbf{A}}\rightarrow{\mathbf{I}}$ be the natural map. The pair $({\mathbf{A}},\nu_{\theta})$ is called an algebra over ${\mathbf{I}}$ in [25, Definition 2.1]. By an old construction of Novotný [28], one can obtain a multisorted algebra ${\mathbf{M}}=\mathfrak{M}({\mathbf{A}},\nu_{\theta})$ whose sort domains are the $\theta$-classes of ${\mathbf{A}}$; then applying another old construction of Gardner [7], one can convert ${\mathbf{M}}$ to a one-sorted algebra whose universe is the Cartesian product of the $\theta$-classes of ${\mathbf{A}}$. Call this final algebra $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$. For precise definitions, see [25, Propositions 2.2 and 2.5]. Because $\theta$ is minimal and abelian, it follows from [25, Corollary 3.5 and Theorem 3.6] that $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ is simple and abelian. And because ${\mathbf{A}}$ has a term $d(x,y,z)$ whose restriction to each $\theta$-class is a Maltsev operation, it follows from the definitions and [25, Remark 2.8] that $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ has a Maltsev term. In fact, writing the universe of $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ as $B_{1}\times\cdots\times B_{k}$ where $B_{1},\ldots,B_{k}$ is an enumeration of the $\theta$-classes, then the Maltsev operation $m(x,y,z)$ of $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ is given coordinatewise by $d$.
@@
Since $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ is simple, abelian, and has a Maltsev term, it follows that $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ is polynomially equivalent to a (finite) simple module by a theorem of Gumm [8] and Herrmann [12]. If we fix a $k$-tuple $\mathbf{e}=(e_{1},\ldots,e_{k})\in B_{1}\times\cdots\times B_{k}$, then we can assume that $\mathbf{e}$ is the zero element of this module and addition is given by $x+y=m(x,\mathbf{e},y)$. By Schur’s Lemma and finiteness, the endomorphism ring of this module is a finite field $\mathbb{F}$, say of cardinality $p^{m}$. It follows that the abelian group $(B_{1}\times\cdots\times B_{k},+,\mathbf{e})$ inherits the structure of a vector space over $\mathbb{F}$, which forces $px=\mathbf{e}$ for all $x\in B_{1}\times\cdots\times B_{k}$. Interpreting this equation coordinatewise gives that each group $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{i})$ satisfies $px=e_{i}$ for each $i=1,\ldots,k$ as required. ∎

### v2

@@ 1. Introduction
Arguably the most important result in universal algebra in the last ten years is the positive resolution of the Constraint Satisfaction Problem (CSP) Dichotomy Conjecture, announced independently in 2017 by Andrei Bulatov [2] and Dmitriy Zhuk [34, 35]. One particular feature of Zhuk’s proof is his analysis of multi-sorted “rectangular critical” relations $R\leq_{sd}{\mathbf{A}}_{1}\times\cdots\times{\mathbf{A}}_{n}$ of finite algebras ${\mathbf{A}}_{1},\ldots,{\mathbf{A}}_{n}$ in certain locally finite idempotent Taylor varieties. Zhuk showed that, in the varieties he considers, such relations produce derived relations, which he called “bridges,” between certain meet-irreducible congruences $\rho_{i}\in\operatorname{Con}({\mathbf{A}}_{i})$ determined by $R$. Zhuk established a number of useful properties of his bridges and used them to tease out implicit linear equations in the CSP instances he considered.
@@
Strikingly, Zhuk’s definition of “bridge” formally resembles some aspects of Hagemann and Herrmann’s formulation [10] from 1979 of the centrality relation on congruences of a single algebra. Moreover, the “there-exists-a-bridge” relation, under suitable restrictions, has some resemblance to the “similarity” relation of Freese [5] from 1983; this latter relation is between subdirectly irreducible algebras in a congruence modular variety, and can be viewed as an amplification of a certain correspondence involving the centrality relation. As well, Zhuk’s bridges arising in rectangular critical subdirect products in locally finite idempotent Taylor varieties resemble the “graphs of similarity” which Kearnes and Szendrei proved arise in rectangular critical relations of the form $R\leq{\mathbf{A}}^{n}$ in congruence modular varieties [19, Theorem 2.5(6)].
@@
Our main goal in this paper and two companion papers [31, 30] is to establish precise connections between Zhuk’s bridges, centrality, and similarity, and then to apply these connections to the study of rectangular critical relations among finite algebras in arbitrary locally finite Taylor varieties. In order to connect Zhuk’s bridges to similarity, we need to first generalize the similarity relation from its original setting (congruence modular varieties) to a broader class also containing all locally finite idempotent Taylor varieties. For this purpose, we focus on the class of varieties which have a weak difference term. Roughly speaking, a “weak difference term” is an idempotent ternary term that specializes to a Maltsev operation whenever it is restricted to a block of an abelian congruence. Such terms are weak versions of terms, known as “difference terms,” that exist and play an important role in congruence modular varieties [9]. Weak difference terms arose organically in tame congruence theory in the 1980s, when Hobby and McKenzie [12] showed that every locally finite Taylor has such a term. The property of a (general) variety having a weak difference term was first explicitly identified in the 1990s in work of Lipparini [21], Kearnes [14], and Kearnes and Szendrei [18], who explored the extent to which centralizer properties in congruence modular varieties extend to varieties with a weak difference term. Recent work concerning varieties with a weak difference term is represented by [16, 33, 20, 15].
This paper is organized around a construction, due to Freese [5] (building on Hagemann and Herrmann [10]) in the congruence modular setting, which we call the difference algebra of an abelian congruence. Whenever $\theta$ is an abelian congruence of an algebra in a variety with a weak difference term, every $\theta$-class naturally supports the structure of an abelian group. Letting $(0:\theta)$ denote the largest congruence centralizing $\theta$, it turns out that the family $\mathscr{F}$ of abelian groups supported on the $\theta$-classes within a fixed $(0:\theta)$-class $E$ can be coherently superimposed on each other, forming a single abelian group $G_{E}$ for which $\mathscr{F}$ is a directed family of subgroups whose union is $G_{E}$. Roughly speaking, the difference algebra of $\theta$ replaces all the $\theta$-classes within each $(0:\theta)$-class $E$ with a single class (of a new abelian congruence $\varphi$) supporting the group $G_{E}$. We define the difference algebra and its congruence $\varphi$, and prove their basic properties, in Sections 4. We make precise the connection between the abelian groups on $\theta$-classes and $\varphi$-classes in Section 5.
@@
In Section 7, we show that Freese’s similarity relation generalizes nicely to varieties with a weak difference term. In the abelian monolith case, Freese’s original definition of similarity, as well as the reformulation in [6], does not generalize in a straightforward way. Instead, we take as our starting point two equivalent (in the congruence modular setting) characterizations of similarity due to Freese – one based on the difference algebra construction, the other a congruence lattice-theoretic formulation – and show that both characterizations generalize to, and are provably equivalent in, varieties with a weak difference term (Theorems 4.6 and 7.7). We also give a third equivalent characterization using a variant of Zhuk’s bridges, which we call “similarity bridges” (Definition 7.3).
@@
Along the way, we expend some energy explaining a folklore result: that each class of an abelian minimal congruence $\theta$ (in an algebra $\mathbf{A}$} with a weak difference term) naturally supports the structure of a left vector space over a division ring $\mathbb{F}_{\theta}$. Here $\mathbb{F}_{\theta}$ depends on $\mathbf{A}$} and $\theta$ but not on the particular $\theta$-class. We provide a new, simpler construction of $\mathbb{F}_{\theta}$ based on the difference algebra construction in Section 6, and show in Appendix 1 that it agrees with the folklore construction. We also show that if two subdirectly irreducible algebras in a variety with a weak difference term have abelian monoliths and are similar, then the division rings determined by their monoliths are isomorphic.
@@
In the companion paper [31], we will develop Zhuk’s theory of bridges and describe the precise relationships between the objects in Zhuk’s theory, the centrality relation, similarity, and similarity bridges as defined here. In a third paper [30], we will study rectangular critical relations in locally finite Taylor varieties, extending the work of Zhuk [35] and connecting these relations to similarity (as was done by Kearnes and Szendrei [19] in the congruence modular setting). In particular, we will show explicitly how each such relations can be understood as encoding a family of “parallel” linear equations over a finite field.
Acknowledgement. I thank Keith Kearnes for several helpful conversations, and Dmitriy Zhuk for providing the inspiration.
@@ 2. Definitions
If $\mathbf{A}$} is an algebra and $\alpha,\beta\in\operatorname{Con}({\mathbf{A}})$ with $\alpha\leq\beta$, then $\beta/\alpha$ denotes the congruence of ${\mathbf{A}}/\alpha$ corresponding to $\beta$ via the Correspondence Theorem ([24, Theorem 4.12] or [1, Theorem 3.6]). If in addition $\gamma,\delta\in\operatorname{Con}({\mathbf{A}})$ with $\gamma\leq\delta$, then we write $(\alpha,\beta)\nearrow(\gamma,\delta)$ and say that $(\alpha,\beta)$ is perspective up to $(\gamma,\delta)$, or transposes up to $(\gamma,\delta)$, if $\beta\wedge\gamma=\alpha$ and $\beta\vee\gamma=\delta$. The notation $(\gamma,\delta)\searrow(\alpha,\beta)$ means the same thing. We write $\alpha\prec\beta$ and say that $\beta$ covers $\alpha$, or is an upper cover of $\alpha$, if $\alpha<\beta$ and there does not exist a congruence $\gamma$ satisfying $\alpha<\gamma<\beta$. A congruence $\alpha$ is minimal if it covers 0, and is completely meet-irreducible if $\alpha\neq 1$ and there exists $\alpha^{+}$ with $\alpha\prec\alpha^{+}$ and such that $\alpha<\beta\implies\alpha^{+}\leq\beta$ for all $\beta\in\operatorname{Con}({\mathbf{A}})$. A subset $T\subseteq A$ is a transversal for a congruence $\alpha$ if it contains exactly one element from each $\alpha$-class.
@@
More generally, if $S,T$ are tolerances of $\mathbf{A}$} (i.e., reflexive symmetric subuniverses of ${\mathbf{A}}^{2}$), then we can define ${\mathbf{A}}(S,T)$, $X(S,T)$, $M(S,T)$, and the relation $C(S,T;\delta)$ in the same way; see [16, §2.5].
Let $\theta,\delta$ be congruences of an algebra $\mathbf{A}$}. The annihilator of $\theta$ modulo $\delta$, denoted $(\delta:\theta)$, is the unique largest congruence $\varphi$ for which $C(\varphi,\theta;\delta)$ holds. We say that $\theta$ is abelian if $C(\theta,\theta;0)$ holds; equivalently, if $\theta\leq(0:\theta)$. More generally, if $\theta,\delta\in\operatorname{Con}({\mathbf{A}})$ with $\delta\leq\theta$, then we say that $\theta$ is abelian modulo $\delta$ if any of the following equivalent conditions hold: (i) $\theta/\delta$ is an abelian congruence of ${\mathbf{A}}/\delta$; (ii) $C(\theta,\theta;\delta)$ holds; (iii) $\theta\leq(\delta:\theta)$. The following lemma expresses that annihilators behave nicely under certain quotients and surjective homomorphisms.
@@
At the level of varieties, the property of having a weak difference term can be characterized by the existence of idempotent terms satisfying a certain pattern of identities, i.e., by an idempotent Maltsev condition, as shown by Kearnes and Szendrei [18, Theorem 4.8]. (See Theorem A2.1 for a slight improvement.) In the context of locally finite varieties, the situation is simpler.
@@
This follows by combining [12, Theorem 9.6], [28, Corollary 5.3], [22, Theorem 2.2], and [18, Theorem 4.8]. ∎
Lemmas 5.2 and 5.7 and Corollary 5.8 mention difference terms; these are weak difference terms which moreover satisfy the identity $d(x,x,y)\approx y$. By an old theorem of Gumm [9], every congruence modular variety has a difference term.
@@
if $a_{i}=b_{i}$ for $i=1,2,3$, then $a_{4}=b_{4}$. It is easy an easy exercise to show that if $\theta$ satisfies the two-term condition, then $\theta$ is abelian. The converse implication is also true if $\mathbf{A}$} belongs to a variety having a weak difference term, by [21, Theorem 3.1] or [18, Corollary 4.5]. (See Lemma 3.3 for a slight improvement.)
@@ 3. Tools
The first result is an easily proved generalization of the well-known fact (see e.g. [1, Theorem 4.65(1)]) that if an algebra $\mathbf{A}$} has a Maltsev term, then every reflexive subuniverse of ${\mathbf{A}}^{2}$ is a congruence of $\mathbf{A}$}. (A proof can be easily extracted from the proof of [16, Corollary 6.22].)
@@
See Lipparini’s proof of [21, Theorem 3.1(a)]. His Theorem 3.1(a) assumes that $\mathbf{A}$} belongs to a variety with a difference term, but in fact all that the argument requires is that $\mathbf{A}$} itself have a weak difference term. ∎
@@
Suppose $\mathbf{A}$} is an algebra having a weak difference term $d(x,y,z)$, and $\theta$ is an abelian congruence of $\mathbf{A}$}. Then for all $e\in A$, the structure $(e/\theta,+,e)$ is an abelian group where $x+y:=d(x,e,y)$ for $x,y\in e/\theta$. Moreover, in this group we have x=d(e,x,e)-x=d(e,x,e) and d(x,y,z)=xy+zd(x,y,z)=x-y+z for all $x,y,z\in e/\theta$.
@@
Let $C=e/\theta$. Then $(C,d{\upharpoonright}_{C^{3}})$ is an abelian algebra with Maltsev operation $d{\upharpoonright}_{C^{3}}$. Now apply Gumm [8, Theorem 4.7] or Herrmann’s main theorem from [11] (or see [29, Corollary 2] for a short proof). ∎
@@
Suppose $\mathbf{A}$} is an algebra and $\theta$ is an abelian congruence. If $d(x,y,z)$ and $t(x,y,z)$ are two weak difference terms for $\mathbf{A}$}, then for every $\theta$-class $C$ we have that $d{\upharpoonright}_{C^{3}}=t{\upharpoonright}_{C^{3}}$.
@@
Again this is a well-known consequence of $(C,d{\upharpoonright}_{C^{3}})$ being an abelian Maltsev operation. For $a,b,c\in C$, Proposition 3.4 yields
d(t(a,b,b),t(b,b,b),t(b,b,c))=t(d(a,b,b),d(b,b,b),d(b,b,c))d(t(a,b,b),t(b,b,b),t(b,b,c))=t(d(a,b,b),d(b,b,b),d(b,b,c))
@@
The operation ++ in the previous definition does not depend on the choice of the weak difference term, by Lemma 3.6. The notation $M(\theta,e)$ has been used in [6] and elsewhere for $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$.
The next result is known in the congruence modular setting (e.g., [6, Chapter 5]), and is folklore for varieties with a weak difference term. We sketch a proof showing that only the algebra $\mathbf{A}$} itself needs to have a weak difference term.
Proposition 3.8.
Suppose $\mathbf{A}$} is an algebra having a weak difference term, and $\theta$ is an abelian congruence.
  1. (1)
    Suppose $C_{1},C_{2}$ are $\theta$-classes and $e_{i}\in C_{i}$ for $i=1,2$. If $r\in\operatorname{Pol}_{1}({\mathbf{A}})$ satisfies $r(e_{1})=e_{2}$, then $r{\upharpoonright}_{C_{1}}$ is a group homomorphism from $\operatorname{Grp}_{{\mathbf{A}}}(\theta_{1},e_{1})$ to $\operatorname{Grp}_{{\mathbf{A}}}(\theta_{2},e_{2})$.
  2. (2)
    Suppose $f\in\operatorname{Pol}_{n}({\mathbf{A}})$ and $C_{1},\ldots,C_{n},C$ are $\theta$-classes with $f(C_{1},\ldots,C_{n})\subseteq C$. Fix $\mathbf{e}=(e_{1},\ldots,e_{n})\in C_{1}\times\cdots\times C_{n}$ and $e\in C$. There exist $r_{1},\ldots,r_{n}\in\operatorname{Pol}_{1}({\mathbf{A}})$ with $r_{i}(e_{i})=e$ for $i\in[n]$, such that for all $\mathbf{a}=(a_{1},\ldots,a_{n})\in C_{1}\times\cdots\times C_{n}$,
[…]
Proof sketch.
(1) follows from Proposition 3.4. (2) is proved by induction on $n$. When $n=1$, define $r_{1}(x)=d(f(x),f(e_{1}),e)$, whose restriction to $C_{1}$ is just $f{\upharpoonright}_{C_{1}}(x)-f(e_{1})$ computed in $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$. When $n>1$, define $g\in\operatorname{Pol}_{n-1}({\mathbf{A}})$ by
$g(x_{1},\ldots,x_{n-1})=f(x_{1},\ldots,x_{n-1},e_{n})$
and inductively get $r_{1},\ldots,r_{n-1}\in\operatorname{Pol}_{1}({\mathbf{A}})$ with $r_{t}(e_{t})=e$ for $t\in[n-1]$, such that $g(a_{1},\ldots,a_{n-1})=\sum_{t=1}^{n-1}r_{t}(a_{t})+f(\mathbf{e})$ for all $\mathbf{a}\in C_{1}\times\cdots\times C_{n-1}$. Define $r_{n}(x)=d(f(e_{1},\ldots,e_{{n-1}},x),f(\mathbf{e}),e)$ and note that $r_{n}(e_{n})=e$. Then for all $\mathbf{a}\in C_{1}\times\cdots\times C_{n}$, computing in $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$,
$\sum_{t=1}^{n}r_{t}(a_{t})+f(\mathbf{e})$ $=\left(\sum_{t=1}^{n-1}r_{t}(a_{t})+f(\mathbf{e})\right)+r_{n}(a_{n})$
$=f(a_{1},\ldots,a_{n-1},e_{n})+(f(e_{1},\ldots,e_{n},a_{n})-f(\mathbf{e}))$
$=d(f(a_{1},\ldots,a_{n-1},e_{n}),f(e_{1},\ldots,e_{n}),f(e_{1},\ldots,e_{{n-1}},a_{n})).$
Using Proposition 3.4, the last expression displayed above is equal to
$f(d(a_{1},e_{1},e_{1}),\ldots,d(a_{n-1},e_{{n-1}},e_{{n-1}}),d(e_{n},e_{n},a_{n}))$
which simplifies to $f(\mathbf{a})$. ∎
The following result is folklore; we include it here for completeness.
@@
This was proved by Freese [5, §4] in the congruence modular case, and essentially the same proof works here. We recast his argument using tools from a recent paper of Mayr and Szendrei [23]. (See also the arguments in Section 6, especially Corollary 6.7, where a generalization is proved.)
@@
First note that we can assume that $\mathbf{A}$} has no nullary operations (by replacing each nullary operation with a constant unary operation). Let ${\mathbf{I}}={\mathbf{A}}/\theta$, and let $\nu_{\theta}:{\mathbf{A}}\rightarrow{\mathbf{I}}$ be the natural map. The pair $({\mathbf{A}},\nu_{\theta})$ is called an algebra over $\mathbf{I}$} in [23, Definition 2.1]. By an old construction of Novotný [26], one can obtain a multisorted algebra ${\mathbf{M}}=\mathfrak{M}({\mathbf{A}},\nu_{\theta})$ whose sort domains are the $\theta$-classes of $\mathbf{A}$}; then applying another old construction of Gardner [7], one can convert $\mathbf{M}$} to a one-sorted algebra whose universe is the Cartesian product of the $\theta$-classes of $\mathbf{A}$}. Call this final algebra $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$. For precise definitions, see [23, Propositions 2.2 and 2.5]. Because $\theta$ is minimal and abelian, it follows from [23, Corollary 3.5 and Theorem 3.6] that $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ is simple and abelian. And because $\mathbf{A}$} has a term $d(x,y,z)$ whose restriction to each $\theta$-class is a Maltsev operation, it follows from the definitions and [23, Remark 2.8] that $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ has a Maltsev term. In fact, writing the universe of $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ as $C_{1}\times\cdots\times C_{k}$ where $C_{1},\ldots,C_{k}$ is an enumeration of the $\theta$-classes, then the Maltsev operation $m(x,y,z)$ of $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ is given coordinatewise by $d$.
@@
Since $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ is simple, abelian, and has a Maltsev term, it follows that $\mathfrak{C}({\mathbf{A}},\nu_{\theta})$ is polynomially equivalent to a (finite) simple module by a theorem of Gumm [8] and Herrmann [11]. If we fix a $k$-tuple $\mathbf{e}=(e_{1},\ldots,e_{k})\in C_{1}\times\cdots\times C_{k}$, then we can assume that $\mathbf{e}$ is the zero element of this module and addition is given by $x+y=m(x,\mathbf{e},y)$. By Schur’s Lemma and finiteness, the endomorphism ring of this module is a finite field $\mathbb{F}$, say of cardinality $p^{m}$. It follows that the abelian group $(C_{1}\times\cdots\times C_{k},+,\mathbf{e})$ inherits the structure of a vector space over $\mathbb{F}$, which forces $px=\mathbf{e}$ for all $x\in C_{1}\times\cdots\times C_{k}$. Interpreting this equation coordinatewise gives that each group $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{i})$ satisfies $px=e_{i}$ as required. ∎